# Uncertainty Quantification of Surrogate Models using Conformal Prediction

### v2

@@ 2.1.2 Definition
Equation 2 is statistically guaranteed to hold for each cell of the spatio-temporal tensor, provided we provide sufficient samples and exchangeability is maintained (Vovk, 2012).
@@ 2.2 Nonconformity Scores
A nonconformity score can be described as a measure of the model’s performance with respect to the calibration dataset. Thus, the scores are a function of the trained model, the calibration inputs, and outputs (Angelopoulos and Bates, 2023). Nonconformity scores describe the deviation of the model from ground truth. In this paper, we focus on three methods of estimating the nonconformity scores for a model:
  • Quantile Regression (CQR): Outlined by (Romano et al., 2019), three models are trained to perform conformal prediction. The lower and upper coverage bands are estimated by models trained to output the $100\times\alpha^{th}$ and $100\times(1-\alpha)^{th}$ percentile, while the median is learned by training the model to output the $50^{th}$ percentile. During training the models are estimated to output the $\alpha^{th}$ percentile by optimising the quantile loss (Koenker, 2005). The nonconformity score measures the distance of the calibration data point from the nearest coverage band: $s(x,y)=\max\{\underline{f}(x)-y,y-\overline{f}(x)\}$. Using the upper and lower quantiles, CQR offers an initial coverage that can be adjusted and calibrated to perform conformal prediction: Using the required quantile ($\hat{q}$), the prediction set is obtained as $\{\underline{f}(x)-\hat{q},\overline{f}(x)+\hat{q}\}$.
  • Absolute Error Residual (AER): Requiring only a single model, trained to approximate the deterministic function itself, this method uses a more straightforward estimate of the model error to calibrate the coverage. nonconformity scores are estimated by taking the absolute error of the model across the labelled calibration dataset (Lei et al., 2018): $s(x,y)=|y-\tilde{f}(x)|$. Upon deriving the quantile ($\hat{q}$), the prediction set is obtained as : $\{\tilde{f}(x)-\hat{q},\;\tilde{f}(x)+\hat{q}\}$. This method requires no structural changes to the model, and only one deterministic model is required to obtain valid prediction sets. The method uses the distribution of the absolute error residuals (AER) from the model to construct sets with the correct coverage.
  • Standard Deviation (STD): The models are probabilistic, i.e., they predict a distribution rather than a deterministic quantity, where the outputs are characterised by a mean $\mu(x)$ and standard deviation $\sigma(x)$. The nonconformity score, $s(x,y):\frac{|y-\mu(x)|}{\sigma(x)}$, takes into account the uncertainty of the model and conformalises it to provide validity over the predictive uncertainty  [^1] If the choice of the model is not probabilistic in itself, we will have to modify the model with changes to architecture or introduce sampling strategies to output a distribution.. The standard deviation of the output distribution provides estimates of the error bars and; thus, it is an initial measure of the coverage, which can be adjusted to ensure the correct coverage using conformal prediction. Upon deriving the quantile ($\hat{q}$), the prediction set is obtained as : $\{\mu(x)-\hat{q}\sigma(x),\;\mu(x)+\hat{q}\sigma(x)\}$. To do CP using a probabilistic model, we modify our networks to behave in a probabilistic manner through architectural modifications or changing the training regime.
Figure 3: Validation plots indicating the coverage guarantee detailed in Equation 1 obtained by CP across experiments, models, and calibration methods. The average empirical coverage guaranteed by CP is given on the $y$-axis (ranging from 0 to 1, with 1 representing 100% coverage), while the coverage, which we calibrate for, is represented on the $x$-axis. Validation plots are shown for the four PDEs we focus on in this paper. We obtain guaranteed coverage for all calibration methods in all our experiments.
Figure 3 illustrates how these nonconformity scores can help obtain guaranteed coverage, across all modelling scenarios considered in the paper(more details on them in Section 3). The coverage provided across the nonconformity scores may vary, but the CP formulation ensures the validity of the error bars. Irrespective of the choice of the nonconformity score, we obtain near-to-ideal coverage for the prediction following the inequality found within Equation 1. Since coverage is guaranteed in the framework, the choice of the nonconformity score is often undertaken considering other aspects such as architectural modifications, calibration time, and cost of obtaining additional data.
The coverage can be empirically tested using a validation dataset. Initially, we perform calibration of the model, which is followed by computing the prediction set $\mathbb{C}^{\alpha}$ using the calibrated metrics, and eventually computing the empirical coverage
@@
where $I_{\mathbb{C}^{\alpha}}$ is the indicator function of the prediction set associated to $\alpha$ prediction set. If the estimated probability is greater than $1-\alpha$, the $\alpha$ prediction sets correctly capture the true coverage.
(Vovk, 2012) shows that the empirical coverage is distributed as
@@
where $l=\lfloor(n_{\text{cal}+1})(1-\alpha)\rfloor$.
CP provides prediction sets at any $\alpha$-level, which coverage may be empirically computed for. This is visualised in Figure 3, where the empirical coverage is plotted against the $1-\alpha$ target coverage from CP. The straight line shows the ideal scenario where the empirical coverage exactly matches the predicted coverage, i.e., the inequality in Equation 1 is an equality. Figure 3 serves as a validation plot of our “conformalised” surrogate model, where we use an unseen validation dataset to empirically check the true coverage (using Equation 3) against the predicted coverage for CP, validating Equation 1 for each $\alpha$-level. Figure 3 shows that we obtain guaranteed coverage across a previously unseen prediction set irrespective of the choice of calibration method across all our experiments.
2.3 Exchangeablity Assumptions
Throughout the experiments considered in this paper, surrogate modelling of spatio-temporal data is treated as an initial value problem (IVP). To map the evolution of the system under study, the surrogate model takes in the initial state(s) of the system and evolves the system autoregressively in time. By treating the modelling task as an initial value problem, each input-output pair ($X_{i},Y_{i}$) from the calibration and the prediction dataset is assumed to be exchangeable. The initial condition associated with each IVP is sampled i.i.d. from a distribution characterising the possible (large) range of interested initial conditions for the prediction task. The assumption relies on the further caveat that is outlined in Section 2.1.2, the structure of each spatio-temporal output is maintained. The CP framework requires exchangeability not just across the calibration data but across the prediction data as well (Angelopoulos and Bates, 2023). To maintain exchangeability across the calibration and prediction datasets, we have to ensure that the full spatio-temporal context window across them remains the same and assume that they arise from the same distribution. The exchangeability assumption is more complex in the case of experimental data, and we discuss this further in Section 3.7 and Section 3.8.
Figure 4: Exchangeable input-output pairs: Within the surrogate models for simulation, the calibration dataset is composed of several simulations, run with different initial conditions, forecastign across the same time domain. Within the surrogate models used for experimental data, exchangeable input-output pairs are formed from within a single experiment, but the IC is taken at several points within the experiment, forecasting across the same time domain.
In Section 3 we consider two types of experiments: surrogate modelling over simulations and experimental data. Within the simulation context (Section 3.1 to Section 3.6), the modelling task takes in the initial state of the system at $t=0$ and then evolves the system (autoregressively or in one-shot) to time $t=T$. The model always starts at the initial condition(s) of the simulation and maps till a fixed time state $T$. Whereas in those modelling experimental data (Section 3.7 and Section 3.8) the modelling task takes the system at $t=T$ to $t=T+\Delta T$. The starting time instance (initial condition) $T$ changes in each exchangeable data pair but the model evolves it for a fixed time duration of $\Delta T$. Within the simulation setup, it is easier to gather data as required, exploring a data-abundant scenario. In the case of real-world experiments, data is scarce and we had to rely on breaking down available data to obtain the calibration data. We characterise this difference in how the exchangeable pairs are conceived for both the simulation and experimental data in Figure 4.
@@ 3 Experiments
To evolve spatio-temporal data describing physical systems, we train several surrogate models using simulation data built from numerical solvers designed for the PDE of interest. Across our experiments, we choose different kinds of surrogate models. Multi-Layer Perceptrons (MLPs) (Haykin, 1994), U-Nets (Ronneberger et al., 2015), Fourier Neural Operators (FNOs) (Li et al., 2021), vision transformers (Yin et al., 2022b), and graph neural networks Scarselli et al. (2009) are utilised. These networks are commonly used as surrogate models for challenging systems in physics and engineering (Mánek et al., 2023; Gupta and Brandstetter, 2023; Wen et al., 2023; Geneva and Zabaras, 2022; Sanchez-Gonzalez et al., 2020).
MLPs have been used as surrogates for modelling wind turbine blades (Lalonde et al., 2021), classifiers for high-energy physics (Baldi et al., 2016), and for designing fusion reactors (Mánek et al., 2023). U-Nets and FNOs are structured as neural-PDE based approximators that can learn across the operator space rather than the function space of the mapping. They find utility in designing surrogates for fluid dynamics (Gupta and Brandstetter, 2023), studying carbon capture and mitigation (Wen et al., 2023), weather modelling (Kurth et al., 2023) and modelling plasma evolution within a nuclear fusion device (Gopakumar et al., 2023). Transformers have found application in PDE solving (Li et al., 2023b) and weather modelling (Nguyen et al., 2023). GNNs have been deployed to model Lagrangian dynamics of a PDE and in solving particle-in-cell scenarios (Brandstetter et al., 2022); they also have been applied to weather forecasting (Lam et al., 2023).
The MLP represents a fully connected neural network. The U-Net is a fully convolutional autoencoder (Hinton and Salakhutdinov, 2006) with skip connections (He et al., 2016; Takamoto et al., 2022). The FNO is an operator learning method that learns the kernel integration across the Fourier space (Li et al., 2021). A vision transformer is an attention-based model that learns spatio-temporal rollouts through sequence patching of the spatial outputs and auto-regressive rollouts. Graph neural networks deploy message-passing across nodes representing the spatio-temporal information to model the evolution of a physical system. Depending on the specific experiment we adjust the model setup to either do a one-step forward-in-time mapping or perform an auto-regressive time roll-out over multiple time steps. All models were trained on an NVIDIA A100 GPU, while the calibration and estimation of prediction sets were done on a standard laptop.
Table 1: Coverage comparison of all models ($\alpha=0.1$). ‘Uncalibrated’ refers to coverage before performing CP (not available for AER method), and ‘calibrated’ is coverage after CP. Across our experiments, models, and calibration methods, we obtain nearly perfect coverage. Tightness is measured as the average width of the error bars across each cell of the spatio-temporal tensor in normalised units. represents experiments where the training set and calibration set are from a different distribution. ‘PT’ refers to the evaluation on the pre-trained model, and ‘FT’ refers to those over the fine-tuned models.
[…]
@@
2D Weather GNN 20,602,232 AER - 91.09 229.23 1.13
(Limited Area) GNN STD 74.62 91.08 309.55 0.96
@@
Table 1 forms the centrepiece of our work. The table shows the coverage obtained when predictions are calibrated for $\alpha=0.1$ (90% coverage) across our experiments utilising the three different calibration methods (nonconformity scores) discussed in Section 2.2. For CQR and STD, we additionally show the coverage before performing CP, as these methods have initial uncertainty estimates, which are calibrated to ensure guaranteed coverage. This is labelled as ‘uncalibrated’ in Table 1. For each experiment, an adequate pre-trained surrogate model is developed or chosen. The CP framework is performed over its output dimensions, observing near-to-ideal coverage. The “tightness of fit” depends on the choice of nonconformity score. The tightness is taken to be the width of the calibrated error bars, averaged across each cell of the tensor. In most experiments we find the AER to offer the best fit, however, STD is a close second.
Demonstrations of the CP formulation for estimating marginal coverage as given in Section 2.1.2 for each experiment, irrespective of the complexity and case-setting follow the structure outlined below.
CP Framework: Experiment Structure
1:Generate training data (if training is involved)
2:Pre-train / fine-tune surrogate model
3:Generate calibration dataset / Use fine-tune dataset
4:Obtain error bars using the Conformal Prediction framework
5:Add the error bars to the prediction datasets
6:Evaluate coverage
@@
The Poisson equation is a generalisation of the Laplace equation. It relates the Laplacian of a field to the distribution of the sources/sinks in the field (Hackbusch, 2017). The elliptic PDE models various physical phenomena, such as electrostatics, gravitation, and fluid dynamics. The Poisson equation is a steady-state problem (spatial-only), where the interest lies in mapping the initial distribution of the field to its final state. The state is given along the x-axis, discretising the domain $[0,1]$ into 32 uniform grid points. An initial dataset consisting of 5000 simulation data points is generated and used to train an MLP. Another 1000 data points are generated to calibrate the trained model using the CP framework along with another 1000 for validating the procedure. The training data, calibration, and validation datasets are sampled from the same distribution within this case. Further details about the physics, data generation strategies and model training can be found in Appendix A.
Figure 5: Error bars obtained by performing CP over the MLP modelling the Poisson equation for the three different methods. Considering the simple setup of the problem, the model fits to near perfection, leading to tighter fits, with small error bars characterising the calibrated bounds.
Multiple MLPs are trained to demonstrate the efficacy of different nonconformity scores while using the CP framework. To perform the CQR, three models are trained, modelling the $5^{th}$, $50^{th}$ and $95^{th}$ quantile, while for AER a single model is trained to output the evolved field using an L1-norm. For CP using STD, an MLP with dropout is trained to obtain a probabilistic model that learns the distribution of the evolved state. The performance comparison of each nonconformity score is laid out in Table 1.
Figure 5 visualises the $\alpha=0.1$ prediction sets ($90\%$ confidence intervals) of CQR (left), AER (centre), and STD (right). Considering the simplicity of the physical dynamics, the MLP learns the mapping to near perfection leading to tighter uncertainty estimates, with a narrow set characterising the uncertainty. [^2] A fully reproducible script demonstrating the CP framework over the Poisson equation can be found at https://github.com/gitvicky/Spatio-Temporal-CP/blob/main/Poisson/Poisson1d_NN_CP.py.
@@
Moving further up in complexity from a steady-state (spatial-only) to a dynamical case (spatio-temporal), consider the convection-diffusion equation in one dimension. As a combination of a parabolic and hyperbolic PDE, the equation is crucial in modelling transport phenomena in various science and engineering applications (Chandrasekhar, 1943). Following the schema laid out in Algorithm CP Framework: Experiment Structure: CP Structure, an initial training dataset of 3000 simulations is generated by sampling the initial conditions and the physical factors of the PDE (diffusion coefficient and convection velocity) from a hypercube. A 1D U-Net is trained as the surrogate to model the spatio-temporal evolution given by the convection-diffusion equation. The model takes in the first 10 time instances to map to the next 10 time instances. The physics, sampling strategies, and model details can be found in Appendix B. The calibration and validation datasets are generated from a slightly shifted domain space as that used for training, where the simulation physics is modelled with less diffusion and more convection. This helps to explore a scenario of deploying a pre-trained surrogate model on an out-of-training-distribution setting.
Figure 6: Calibrated lower and upper bounds of the model output for the 1D convection-diffusion equation with $\alpha=0.1$ providing 90% coverage. The plots show the field distribution across the $x$-axis at the final time instance $t=0.1$. When the uncalibrated coverage is too constrained (as with CQR), CP helps widen it to provide the required coverage. In the case of STD, note that for models with lower dimensionality, the sampled uncertainty is calibrated moderately well before performing CP over it.
@@
The wave equation is a second-order hyperbolic partial differential equation that describes the propagation of waves through a medium with numerous applications in acoustics, optics and quantum mechanics (Tipler, 2008). The experiment focuses on the spatio-temporal dynamics of a 2D Gaussian under the constraints of the wave equation. The detailed description of the physics, the numerical solvers, and the data generation strategy are given in Appendix C. To build the pre-trained model to evaluate the CP framework, initially, a training dataset of 500 simulations is generated by varying the parameterisation of the initial condition, i.e., the amplitude and position of the Gaussian. The data is used to train a U-Net and an FNO that models the temporal evolution of the wave dynamics.

### v3

@@ 2.1.1 Mathematical Formulation
Equation 2 holds for each tensor cell given sufficient calibration samples and maintained exchangeability (vovk2012conditional).
@@ 2.2 Nonconformity Scores
Nonconformity scores quantify model deviation from ground truth using the calibration set (gentle_introduction_CP). We employ three methods:
  • Conformalised Quantile Regression (CQR) (conformalized_quantile_regression): Train three models to predict the $100\times\alpha$-th, median, and $100(1-\alpha)$-th percentiles using quantile loss (koenker_2005). The nonconformity score measures distance to the nearest bound: $s(x,y)=\max\{\underline{f}(x)-y,y-\overline{f}(x)\}$. After computing $\hat{q}$, the prediction set is obtained as $\{\underline{f}(x)-\hat{q},\overline{f}(x)+\hat{q}\}$.
  • Absolute Error Residual (AER) (error_residual): Train a single deterministic model using standard loss (e.g., MSE). Compute nonconformity scores as absolute errors: $s(x,y)=|y-\tilde{f}(x)|$. The prediction set becomes $\{\tilde{f}(x)-\hat{q},\tilde{f}(x)+\hat{q}\}$. This requires no architectural modifications and is computationally efficient.
  • Standard Deviation (STD): Use probabilistic models outputting mean $\mu(x)$ and standard deviation $\sigma(x)$. The nonconformity score is $s(x,y)=\frac{|y-\mu(x)|}{\sigma(x)}$, yielding prediction sets $\{\mu(x)-\hat{q}\sigma(x),\mu(x)+\hat{q}\sigma(x)\}$. This requires architectural changes (e.g., dropout layers) or modified training (e.g., negative log-likelihood loss). The dependence on the standard deviation of the prediction introduces a weak sense of conditionality.
Figure 3: Empirical coverage versus target coverage $(1-\alpha)$ across experiments and nonconformity scores. The diagonal represents ideal coverage. All methods achieve near-perfect coverage across four PDE experiments, validating Equation 1.
Figure 3 demonstrates that all three nonconformity scores achieve guaranteed coverage across diverse experiments (details in Section 3). While coverage quality varies slightly between methods, CP ensures validity regardless of the choice. We select nonconformity scores based on practical considerations: architectural constraints, calibration cost, and data availability.
To validate coverage empirically, we compute:
@@
where $I_{\mathbb{C}^{\alpha}}$ is the indicator function for the prediction set. Valid coverage requires this to exceed $1-\alpha$. The empirical coverage obtained for a setting follows a Beta distribution characterised as (vovk2012conditional):
@@
where $l=\lfloor(n_{\text{cal}+1})(1-\alpha)\rfloor$ and $n_{\text{cal}}$ is the calibration set size.
2.3 Computational Complexity of Calibration
The conformal prediction calibration procedure has complexity $\mathcal{O}(d\cdot n_{\text{cal}}\log n_{\text{cal}})$, where $d=T_{\text{out}}\times N_{x}\times N_{y}\times N_{\text{var}}$ is the output dimensionality and $n_{\text{cal}}$ is the calibration set size. This comprises: (1) computing nonconformity scores $s(x,y)$ for all $n_{\text{cal}}$ samples across $d$ dimensions, requiring $\mathcal{O}(n_{\text{cal}}\cdot d)$ element-wise operations (e.g., |YY~||Y-$\tilde${Y}| for AER), and (2) sorting scores to estimate quantiles $\hat{q}$ per dimension, requiring $\mathcal{O}(d\cdot n_{\text{cal}}\log n_{\text{cal}})$ operations. Once calibrated, constructing prediction sets for new predictions requires only $\mathcal{O}(d)$ operations by applying the pre-computed quantiles. The procedure requires only forward passes and sorting—no gradient computation or model retraining—and is embarrassingly parallel across dimensions. This computational efficiency is particularly advantageous compared to alternative UQ methods that require ensemble training ($\mathcal{O}(n_{\text{ensemble}}\cdot\text{training cost})$) or extensive Bayesian sampling. As shown in Table 1, calibration remains practical even for $d>10^{7}$ dimensions, with times ranging from <1<1 second (low-dimensional cases) to a few hundred seconds (20M+ dimensions for weather forecasting), demonstrating the near-zero computational cost that makes CP particularly suitable for production deployment of surrogate models.
2.4 Exchangeability Requirements
We treat spatio-temporal surrogate modelling as an initial-value problem (IVP), where models evolve initial states autoregressively or in one-shot mappings. Each input-output pair $(X_{i},Y_{i})$ from calibration and prediction sets is assumed exchangeable when initial conditions are sampled i.i.d. from the distribution of interest. This assumption requires: (1) consistent spatio-temporal structure (Section 2.1.1) across the inputs and output, (2) identical discretised domains across calibration and prediction, and (3) exchangeability between calibration and test data (gentle_introduction_CP).
Figure 4: Constructing exchangeable pairs for simulation versus experimental data. Simulation-based: Multiple independent runs with varying initial conditions at $t=0$, each forecasting the same time horizon to $t=T$. Experimental data: Single long-duration experiment partitioned into multiple IVPs with different starting times $T_{i}$ from a larger temporal domain, each forecasting the same duration $\Delta T$.
Our experiments (Section 3) involve two distinct scenarios for constructing exchangeable pairs (Figure 4):
2.4.1 Simulation-Based Exchangeability
For surrogate models of numerical simulations (Section 3.1Section 3.6), exchangeability is straightforward. The modelling task maps from the simulation start $t=0$ to a fixed future time $t=T$. Each simulation begins at $t=0$ with initial conditions sampled from a distribution $\mathcal{P}_{\text{IC}}$ characterising the parameter space of interest. For example, in the MHD plasma blob experiments (Section 3.5), initial conditions vary in blob positions, widths, and amplitudes (Table 9), while the temporal evolution window $[0,T]$ remains identical across all simulations.
This creates a natural exchangeability structure: each pair $(X_{i},Y_{i})$ represents an independent draw from the joint distribution of initial conditions and their corresponding solutions. Multiple simulations with varied physical parameters provide abundant exchangeable pairs without violating temporal dependencies, as each simulation is an independent realisation of the physical system. The computational cost of generating additional calibration data is primarily limited by simulation expense rather than data availability.
2.4.2 Experimental Data Exchangeability
For surrogate models trained on experimental observations (Section 3.7, Section 3.8), exchangeability is constructed through temporal windowing while preserving the IVP structure. Experimental data, such as continuous weather observations or plasma diagnostics, form long time series from which we extract multiple training examples.
The key insight is constructing exchangeable IVPs from a single temporal sequence. Each pair $(X_{i},Y_{i})$ represents an IVP where: (1) initial condition $X_{i}$ is extracted at time $t=T_{i}$ from the larger domain, (2) the model forecasts forward for fixed duration $\Delta T$, (3) target $Y_{i}$ spans $T_{i},T_{i}+\Delta T]$, and (4) starting time $T_{i}$ varies while $\Delta T$ remains constant. Here, these pairs can be treated as exchangeable, as we consider the data distribution to be characterised by a wide range of initial conditions, and the prediction window remains the same.
Crucially, predictions depend only on $X_{i}$, independent of absolute time $T_{i}$. A weather forecast initialised at 12:00 January 5th, predicting 48 hours ahead is exchangeable with one initialised at 18:00 January 12th with the same forecast horizon, such that both solve the same IVP: “given these atmospheric conditions, predict evolution over 48 hours.” This time windowing transforms a single time series into $N$ exchangeable pairs by sampling initial conditions $T_{1},T_{2},\ldots,T_{N}$ from distribution $\mathcal{P}_{T}$ representing typical states within the observed period. The fixed window $\Delta T$ ensures identical structure across pairs, satisfying Section 2.1.1.
For fusion camera diagnostics (Section 3.8), we extract 10-frame initial conditions from different time points within discharge shots, each predicting the subsequent 10 frames. The model learns mappings from the current plasma state to the near-future state, agnostic to absolute discharge time. Similarly, weather forecasting (Section 3.7) constructs pairs from different initialisation times across months, each representing the same multi-day forecast problem.
Important limitation: Exchangeability can be violated with experimental data when calibration and test distributions differ (e.g., different seasons in weather, different plasma regimes in fusion). We demonstrate this sensitivity in Section 3.7 and Section 3.8, showing reduced coverage when exchangeability assumptions break. Users must verify distributional similarity or recalibrate when conditions change substantially.
@@ 3 Experiments
We empirically validate our CP framework across diverse surrogate models trained on spatio-temporal data from physical systems. Our experiments span multiple neural architectures commonly deployed in scientific modelling: Multi-Layer Perceptrons (MLPs) (haykin1994neural), U-Nets (ronneberger2015unet), Fourier Neural Operators (FNOs) (li2021fourier), vision transformers (Yin_2022_CVPR), and graph neural networks (GNNs2009). These architectures have proven effective for surrogate modelling in diverse applications, including wind turbine design (LALONDE2021104696), high-energy physics (Baldi2016), fusion reactors (Manek_2023), fluid dynamics (gupta2022multispatiotemporalscale), carbon capture (Wen_2023), weather forecasting (pathak2022fourcastnet; lam2022graphcast), and plasma evolution (GOPAKUMAR2023100464).
Relationship to Other UQ Methods:
In this work, we focus on empirically demonstrating how inductive CP provides calibrated error bars across diverse spatio-temporal models using various non-conformity scores (as detailed in Section 2.2). Our framework addresses both deterministic models (using AER) and probabilistic models—whether frequentist (CQR) or Bayesian (STD)—showcasing CP’s model-agnostic nature and its ability to provide or calibrate guaranteed coverage across these different paradigms.
CP fundamentally differs from alternative UQ approaches such as deep ensembles and Bayesian neural networks by providing finite-sample, distribution-free guarantees on coverage. In contrast, ensemble and Bayesian methods provide asymptotic or model-dependent uncertainty estimates without such guarantees. While these methods can produce useful uncertainty estimates, they often fail to achieve desired coverage without post-hoc calibration. A comprehensive empirical comparison between CP and Bayesian deep learning methods for surrogate modelling—including coverage reliability, computational costs, and calibration quality—is presented in gopakumar2025calibrated, where we demonstrate that methods like MC Dropout and deep ensembles frequently require further calibration to achieve valid coverage. The present work establishes CP’s applicability and scalability to high-dimensional spatio-temporal problems across multiple scientific domains.
Computational Setup:
All models were trained on NVIDIA A100 GPUs with 80GB memory. Calibration and prediction sets were evaluated on standard laptop hardware, demonstrating the computational efficiency of our approach. The experiment procedure is outlined in the Algorithm CP Framework: Experiment Structure below.
CP Framework: Experiment Structure CP Framework Structure
1:Generate/gather training data (simulation or experimental)
2:Train surrogate model (or use pre-trained model)
3:Generate/gather calibration dataset (or use fine-tuning dataset)
4:Compute nonconformity scores and quantiles via CP framework
5:Construct prediction sets with guaranteed coverage
6:Validate coverage on independent test set
Table 1: Comprehensive coverage results for $\alpha=0.1$ (90% target coverage). Uncalibrated coverage (unavailable for AER) shows initial estimates before CP calibration. Calibration time is reported on standard laptop hardware. Tightness represents average error bar width in normalised units (Min-Max normalisation between -1 and 1). indicates out-of-distribution evaluation where calibration and training distributions differ. PT: pre-trained model; FT: fine-tuned model.
Case Model Output Dims Method Uncalib. (%) Calib. (%) Cal. Time (s) Tightness
@@
2D Weather GNN 20,602,232 AER - 91.19 229.23 1.13
(Limited Area) GNN STD 73.55 89.97 309.55 0.91
@@
Table 1 presents our main empirical findings. Across all experiments, models, and nonconformity scores, we achieve near-perfect calibration to the target 90% coverage ($\alpha=0.1$). For methods with initial uncertainty estimates (CQR and STD), we show both uncalibrated and calibrated coverage, demonstrating how CP corrects potentially misleading uncertainty quantification. The tightness metric, computed as average error bar width in the normalised space (linear transformation of the field to lie between -1 and 1), reveals that AER generally provides the tightest fits, with STD as a close second. Critically, these results hold across output dimensions ranging from 32 to over 20 million, demonstrating that our framework overcomes the curse of dimensionality. All models were trained on an Nvidia A100 GPU and evaluated and calibrated over standard laptop hardware.
@@
The Poisson equation generalises the Laplace equation and models diverse phenomena, including electrostatics, gravitation, and fluid potential fields (Hackbusch2017). This steady-state elliptic PDE serves as our simplest test case, mapping an initial field distribution to its equilibrium state along a 1D domain $[0,1]$ discretised into 32 uniform points.
Dataset: We generated 7,000 simulations using finite difference methods (py-pde package (py-pde)) by varying the initial field value $u_{\text{init}}\sim\mathcal{U}(0,4)$, allocated as 5,000 training, 1,000 calibration, and 1,000 validation samples. All datasets were sampled from the same distribution.
Models and Training: We trained separate MLPs (3 layers, 64 neurons per layer) for each nonconformity method: (i) three models for CQR modeling the $5^{\text{th}}$, $50^{\text{th}}$, and $95^{\text{th}}$ quantiles using quantile loss (koenker_2005); (ii) one deterministic model for AER using L1 loss; and (iii) one probabilistic model with dropout layers for STD. Training used the Adam optimiser (adam) with initial learning rate 0.005 (halved every 100 epochs) for up to 1,000 epochs. Further details about the physics, data generation strategies and model training can be found in Appendix A.
Figure 5: Calibrated prediction sets ($\alpha=0.1$, 90% coverage) for the 1D Poisson equation using three nonconformity scores (CQR, AER, STD). The simple dynamics enable near-perfect model fit, resulting in tight, well-calibrated error bars. Ground truth (black line), model prediction (blue line), and shaded regions depict the prediction sets.
Results: Figure 5 visualizes the $\alpha=0.1$ prediction sets for all three methods. The MLP learns this simple mapping with high accuracy, yielding tight uncertainty bounds. All methods achieve the guaranteed coverage (Table 1), with AER providing the tightest fit (0.002 normalised units) and minimal calibration time (3 ms). The probabilistic STD method achieves comparable coverage but with slightly wider bounds (0.025 normalised units) and longer calibration time (133 ms) due to Monte Carlo dropout sampling.
@@
We advance to a spatio-temporal system governed by the convection-diffusion equation, which combines parabolic and hyperbolic PDE characteristics to model transport phenomena across diverse applications (Chandrasekhar). This equation describes how a fluid density field evolves under the competing effects of diffusion (smoothing) and convection (transport). We consider a spatially-varying diffusion coefficient and Gaussian initial conditions parametrised by mean and variance, with the system evolving over a 1D spatial domain $x\in[0,10]$ and time interval $t\in[0,0.1]$.
Dataset: We generated 5,000 simulations using a forward-time centred-space finite difference scheme across 200 spatial points and 100 time steps. The training set (3,000 samples) was generated via Latin hypercube sampling over physically relevant ranges of diffusion coefficients, convection velocities, and initial condition parameters. To test out-of-distribution robustness, calibration and validation sets (1,000 each) were sampled from a shifted parameter regime with reduced diffusion and enhanced convection. This distribution shift mimics real-world deployment scenarios where test conditions may differ from training.
Model and Training: A 1D U-Net with 4 encoder-decoder levels maps the first 10 time steps (downsampled from 100) to the next 10 steps. The architecture uses batch normalisation and tanh activations. For STD, we added dropout layers (rate 0.1) after each encoder-decoder block. Training followed the same optimiser schedule as the Poisson case, using quantile loss for CQR and MSE loss otherwise. Further details about the physics, data generation strategies and model training can be found in Appendix B.
Figure 6: Calibrated prediction sets ($\alpha=0.1$, 90% coverage) for 1D convection-diffusion evaluated on out-of-distribution data at final time $t=0.1$. CP successfully calibrates initially insufficient coverage (CQR: 25.53% $\rightarrow$ 93.05%) or refines sampled uncertainty (STD: 88.43% $\rightarrow$ 90.29%) to guarantee target coverage. Ground truth (black), prediction (blue), calibrated bounds (red shaded), and uncalibrated bounds (gray dashed) where applicable.
@@
The 2D wave equation models wave propagation in acoustics, optics, and quantum mechanics (thermodynamics_textbook). We simulate Gaussian wave packets evolving on a $33\times 33$ spatial grid over 80 time steps. A training dataset of 500 simulations is generated by varying the amplitude and position of the initial Gaussian. We train both a U-Net (feed-forward) and an FNO (autoregressive) to model the temporal evolution. Calibration and validation each use 100 additional simulations respectively. Full physics details and numerical methods are in Appendix C.